# Model Alignment Search

### v1

@@ Abstract
When can we say that two neural systems are the same? The answer to this question is goal-dependent, and it is often addressed through correlative methods such as Representational Similarity Analysis (RSA) and Centered Kernel Alignment (CKA). What do we miss when we forgo causal explorations, and how can we target specific types of similarity? In this work, we introduce Model Alignment Search (MAS), a method for causally exploring distributed representational similarity. The method learns invertible linear transformations that align a subspace between two distributed networks’ representations where causal information can be freely interchanged. We first show that the method can be used to transfer specific causal variables, such as the number of items in a counting task, between networks with different training seeds. We then explore open questions in number cognition by comparing different types of numeric representations in models trained on structurally different numeric tasks. We then explore differences between MAS vs preexisting causal similarity methods, showing MAS to be more resistant to unwanted exchanges. Lastly, we introduce a counterfactual latent auxiliary loss function that helps shape causally relevant alignments even in cases where we do not have causal access to one of the two models for training.
Model Similarity, DAS, Math Cog, Numbers, Counting
@@ 2 Methods
In this work, we build upon the work of (Grant et al., 2024) to examine the causal similarity of distributed representations within models trained on next token prediction, numeric tasks. Each model is trained to ¿99.99% accuracy on both training and validation data held before being analyzed and interpreted.
@@ 2.4 Model Alignment Search (MAS)
We first train the models to ¿99.99% accuracy, evaluated on both training and held out validation data (we note that MAS can be used on models that are not fully trained). We freeze the model weights for the MAS analysis. MAS can be thought of as a multi-model extension of DAS (Geiger et al., 2021, 2023) where the aligned subspace is used for within model causal interventions (like in DAS) and between model interventions (similar to model stitching). We provide a visual overview in Figure 1. For each distributed system in the analysis, MAS learns an invertible transformation matrix, $\mathcal{R}_{i}=c_{i}Q_{i}$, composed of an orthogonal rotation matrix, $Q_{i}\in R^{m\times m}$, and a scaling constant, $c_{i}$. The transformation matrices rotate the models’ representations into a space where causally important information can be exchanged within and between each of the models included in the analysis. These rotation matrices are trained and evaluated using expected counterfactual behavior as NTP labels.
@@ 3.4 Direct Linear Mappings
How is MAS different than the direct linear mappings that have been used in prior work (Sexton & Love, 2022; Lähner et al., 2023; Lenc & Vedaldi, 2015; Bansal et al., 2021; Sucholutsky et al., 2023)? As far as we are aware, the direct linear mappings in previous work have not isolated specific subspaces for transfer as we have done in this work. Using this subspace isolation, MAS allows the examination of individual causal variables (as opposed to all causal variables) while constraining the alignment to also work within each individual model. A less obvious benefit is in cases where we wish to ignore task irrelevant variability in the target model’s activations when performing an intervention. It can sometimes be unclear how a portion of variability in the activations contributes to a task, but that variability can still be important for the models’ behavior (Li & McClelland, 2022). MAS allows for this type of activity to persist during transfers by selecting only a causal subspace for the transfer.
Turning now to Figure 5, we see three components of a theoretical models’ state vector and their corresponding values at different points along 3 different trajectories. Each trajectory begins at the 0,0,0 coordinate. Each blue arrow represents an increment in the number of demo tokens, allowing the total number of demo tokens to be represented as the magnitude of the z-component. The number of response tokens is nonlinearly encoded along the y-axis, and the red curve represents the model’s decision boundary for outputting the E token. For a single trajectory, this model effectively counts up along the z-axis, then counts along the y-axis, and finishes when it has reached the non-linear decision boundary. Along the x-axis, we show an encoding correlate that does not affect the model’s functional solution to the task, but still encodes relevant information that can be used for neural decoding. In this model, it is possible to linearly decode the Count using the difference between the magnitude of the x-component and the magnitude of the z-component. It is possible for this Count to then be mapped to the z-axis equivalent of a similar target model which could, in turn, continue counting demo tokens or begin counting response tokens for appropriate counterfactual behavior. It is important to note that this substitution is likely only possible if the mapping can push all other causally relevant information into the null space of the target model, as remnants of the z or y components would interfere with the appropriate counterfactual behavior. This effect is one potential explanation for why the IIA observed in MAS is lower; namely, the within model interventions hinder solutions that ”throw away” residual information. We do not make efforts to prove that this theoretical model accounts for the results we see in Figure 4, however, it serves as a grounding for understanding one way that we could observe the results.

### v2

@@ Abstract
When can we say that two neural systems are the same? The answer to this question is goal-dependent, and it is often addressed through correlative methods such as Representational Similarity Analysis (RSA) and Centered Kernel Alignment (CKA). What do we miss when we forgo causal explorations, and how can we target specific types of similarity? In this work, we introduce Model Alignment Search (MAS), a method for causally exploring distributed representational similarity. The method learns invertible linear transformations that align a subspace between two distributed networks’ representations where causal information can be freely interchanged. We first show that the method can be used to transfer specific causal variables, such as the number of items in a counting task, between networks with different training seeds. We then explore open questions in number cognition by comparing different types of numeric representations in models trained on structurally different numeric tasks. We then explore differences between MAS vs preexisting causal similarity methods, and lastly, we introduce a counterfactual latent auxiliary loss function that helps shape causally relevant alignments even in cases where we do not have causal access to one of the two models for training.
Model Similarity, DAS, Math Cog, Numbers, Counting, Arithmetic
@@ 2 Methods
In this work, we build upon the work of (Grant et al., 2024) to examine the causal similarity of distributed representations within models trained on next token prediction, numeric tasks. Each model is trained to $>99.99\%$ accuracy on both training and validation data held before being analyzed and interpreted.
@@ 2.4 Model Alignment Search (MAS)
We first train the models to $>99.99\%$ accuracy, evaluated on both training and held out validation data (we note that MAS can be used on models that are not fully trained). We freeze the model weights for the MAS analysis. MAS can be thought of as a multi-model extension of DAS (Geiger et al., 2021, 2023) where the aligned subspace is used for within model causal interventions (like in DAS) and between model interventions (similar to model stitching). We provide a visual overview in Figure 1. For each distributed system in the analysis, MAS learns an invertible transformation matrix, $\mathcal{R}_{i}=c_{i}Q_{i}$, composed of an orthogonal rotation matrix, $Q_{i}\in R^{m\times m}$, and a scaling constant, $c_{i}$. The transformation matrices rotate the models’ representations into a space where causally important information can be exchanged within and between each of the models included in the analysis. These rotation matrices are trained and evaluated using expected counterfactual behavior as NTP labels.
@@ 3.4 Direct Linear Mappings
How is MAS different than the direct linear mappings that have been used in prior work (Sexton & Love, 2022; Lähner et al., 2023; Lenc & Vedaldi, 2015; Bansal et al., 2021; Sucholutsky et al., 2023)? As far as we are aware, the direct linear mappings in previous work have not isolated specific subspaces for transfer as we have done in this work. Using this subspace isolation, MAS allows the examination of individual causal variables (as opposed to all relevant causal variables) while constraining the alignment to also work within each individual model. A less obvious benefit is in cases where we wish to ignore task irrelevant variability in the target model’s activations when performing an intervention. It can sometimes be unclear how a portion of variability in the activations contributes to a task, but that variability can still be important for the models’ behavior (Li & McClelland, 2022). MAS allows for this type of activity to persist during transfers by selecting only a causal subspace for the transfer.
Turning now to Figure 5, we see three components of a theoretical models’ state vector and their corresponding values at different points along 3 different trajectories. Each trajectory begins at the 0,0,0 coordinate. Each blue arrow represents an increment in the number of demo tokens, allowing the total number of demo tokens to be represented as the magnitude of the z-component. The number of response tokens is encoded nonlinearly along the y-axis, and the red curve represents the model’s decision boundary for outputting the E token. For a single trajectory, this model effectively counts up along the z-axis, then counts along the y-axis, and finishes when it has reached the non-linear decision boundary. Along the x-axis, we show an encoding correlate that does not affect the model’s functional solution to the task, but still encodes relevant information that can be used for decoding. In this model, it is possible to linearly decode the Count using the difference between the magnitude of the z-component and the magnitude of the x-component. More precisely, $\text{Count}=2z-x$. It is possible to then map this information to the z-axis equivalent of a similar target model which could, in turn, continue counting demo tokens or begin counting response tokens for appropriate counterfactual behavior. It is important to note that this substitution is likely only possible if the mapping can push all other causally relevant information into the null space of the target model, as remnants of the z or y components would interfere with the appropriate counterfactual behavior. This effect is one potential explanation for why the IIA observed in MAS is lower. Namely, the within model interventions hinder solutions that ”throw away” residual information. We do not make efforts to prove that this theoretical model accounts for the results we see in Figure 4, however, it serves as one potential way that we could observe our results.